OpenStudy (anonymous):

Graph the function f(x)=1/2[x]

OpenStudy (anonymous):

well look at it like this f(x) = (1/2)x suppose f(x) is called y y = 1/2(x) + 0 so you know the equation of a line is y = mx + b where m is the slope and b is the y intercept

OpenStudy (anonymous):

so would that just be one point on the graph?

OpenStudy (anonymous):

no, that would be a line with its y intercept at (0,0) and with a constant slope of 1/2